How many times do you need to toss a fair coin to get 100 heads with probability 90%?



1
Expert's answer
2022-04-26T00:56:18-0400

Since a fair coin is considered, the probability of getting a head is "\frac12" and the probability of getting a tail is "\frac12". Assume that "X" is a random variable, which denotes a number of heads after "n" tosses. It has a binomial distribution. The aim is to find such "n" that "P(X=100)=0.9." Since "X" has a binomial distribution, we get: "P(X=100)=C_{n}^{100}(\frac12)^{n}", where "C_n^{100}=\frac{n!}{100!(n-100)!}=\frac{(n-99)(n-98)\ldots n}{100!}".

Denote "f(n)=\frac{(n-99)(n-98)\ldots n}{100!}(\frac12)^{n}". The function is considered for "n\geq100". We point out that "f(n)>0". Consider "\frac{f(n+1)}{f(n)}=\frac{(n-98)(n-97)\ldots (n+1)}{(n-99)(n-98)\ldots n}\cdot\frac12=\frac12\cdot\frac{n+1}{n-99}".Consider inequalities: "\frac12\cdot\frac{n+1}{n-99}<1\Longleftrightarrow n+1<2n-198\Longleftrightarrow n>199". Thus, "f(n+1)<f(n)" for "n>199". It means that "f(200)>f(201)>f(202)\ldots"

"\frac12\cdot\frac{n+1}{n-99}>1\Longleftrightarrow n+1>2n-198\Longleftrightarrow n<199". Thus, "f(n+1)>f(n)" for "n<199". It means that "f(199)>f(198)>f(197)>\ldots". Thus, it is enough to compute "f(199)" and "f(200)."

"f(199)\approx0.0563" and "f(200)\approx0.0563". Thus, the probability of getting "100" heads is always less than "6\%" irrespectively of the number of tosses.